NH4+ salt is confirmed8/24 points Previous Answers My Note Calculate the change in pH to 0.01 pH units caused by adding 10. mL of 2.82-M NaOH is ad
Answer #1

b) 0.127 M NHAU Kb of NH3 = 1.8 x 10-5 K 1x10-14 K, of NH K 1.8x10-5 = 5.56x10-10 The dissociation of NH4 is given by К. | NH

Addition of NaOH: Number of moles of NH4+ =M*V = 0.127 M * 390 mL = 49.53 mmol Number of moles of NaOH = M*V = 2.82 M*10 mL =

NH,+: pH before mixing = 5.08 pH after mixing = 9.38 pH change = 9.38 – 5.08 = 4.30 c) 0.127 M NH3 Ko of NH3 = 1.8 x 10-5 The

Addition of NaOH: Number of moles of NH3=M*V = 0.127 M * 390 mL = 49.53 mmol Number of moles of NaOH = M*V = 2.82 M*10 mL = 2

NHz: pH before mixing = 11.18 pH after mixing = 12.85 pH change = 12.85 - 11.18 = 1.67 (D) Solution containing both NHz and N

Addition of NaOH: Number of moles of NH+ =M*V = 0.127 M * 390 mL = 49.53 mmol Number of moles of NH4+ =M*V = 0.127 M * 390 mL

(d) For NH, and NHAT : pH before mixing = 9.26 pH after mixing = 9.82 pH change = 9.82 - 9.26 = 0.56